# Color-avoiding Ramsey for directed paths and vector sequences

### v1

@@ Abstract.
We study two closely related Ramsey-type questions. The first is about vectors: how long can a sequence of vectors in $n]^{q}$ be if, for all pairs in the sequence, the later vector is strictly larger than the earlier in at least q1q-1 coordinates? The second is graph-theoretic: given a $q$-edge-coloring of a tournament, how long of a directed path can we guarantee whose edges avoid one of the colors? These problems arise naturally in many areas, such as convex geometry and extremal hypergraph theory, and have been extensively studied over the past 50 years.
We prove that if $\varepsilon>0$ is fixed and $q$ is sufficiently large, then every $N$-vertex tournament contains a color-avoiding directed path of length $N^{1-\varepsilon}$. Consequently, the maximum length of a (q1)(q-1)-increasing sequence of vectors in $n]^{q}$ has length at most $n^{1+\varepsilon}$. Our results answer a question of Gowers and Long, strengthen several of their results, and extend earlier works of Tiskin and Loh.
JF: Department of Mathematics, Stanford University, USA. Email: jacobfox@stanford.edu. Research supported by NSF Award DMS-2154129
@@ 1. Introduction
One of the earliest, and most influential, results in Ramsey theory is the Erdős–Szekeres lemma [15], which states that every sequence of $N$ real numbers contains a monotone subsequence of length at least $\sqrt{N}$. This innocent statement turns out to have deep connections to discrete geometry [15], high-dimensional partitions [27], Ramsey theory of paths [19], and sequences of vectors [31]; the latter two will be of particular importance to us, as we shortly explain.
The Erdős–Szekeres lemma has an equivalent formulation in terms of edge colorings: it states that in every 2-edge-coloring of the vertices of a complete graph with vertex set $[N]$, there is a monochromatic monotone path of length $\sqrt{N}$; here, and throughout the paper, a path is monotone if its vertices come in increasing order, and the length of a path refers to its number of vertices. A particularly elegant proof of this statement, which explains the connections to relations on vectors, was given by Seidenberg [31]: one simply records for each vertex the length of the longest monotone monochromatic path ending at that vertex in each color, and notes that these vectors must be pairwise distinct, from which the result follows by the pigeonhole principle. More precisely, given two vectors $x,y\in\mathbb{R}^{q}$, let us write that $x<_{1}y$ if there is at least one coordinate $i\in[q]$ such that $x_{i}<y_{i}$; then Seidenberg’s argument shows that the Erdős–Szekeres lemma is equivalent to the statement that if a sequence of vectors $x_{1},\dots,x_{N}\in[n]^{2}$ has the property that $x_{a}<_{1}x_{b}$ for all $a<b$, then $N\leqslant n^{2}$. And this last statement is essentially trivial, since the pigeonhole principle implies that if $N>n^{2}$, then two vectors in the sequence must be equal, which clearly violates this $1$-increasing property.
There are numerous ways of generalizing these results. For example, the $q$-dimensional analogue of the argument above immediately implies that if we color $E(K_{N})$ with $q$ colors, then we must always have a monochromatic monotone path of length at least $N^{1/q}$. For another natural extension, we can look at arbitrary tournaments; note that a monotone path in $K_{N}$ is precisely the same as a directed path in the transitive $N$-vertex tournament, but the same question makes sense in any tournament. In particular, as observed by Chvátal and Gyárfás–Lehel [9, 23] it is not hard to use the Gallai–Hasse–Roy–Vitaver theorem [21, 25, 29, 38] to show that no matter how we $q$-color the edges of an $N$-vertex tournament, there is a monochromatic directed path of length at least $N^{1/q}$. In this setting, the corresponding vector question asks about the maximum size of a set $\{x_{1},\dots,x_{N}\}\subseteq[n]^{q}$ with the property that for all $a\neq b$, either $x_{a}<_{1}x_{b}$ or $x_{b}<_{1}x_{a}$; that is, we still want every pair to be strictly comparable in at least one coordinate, but we do not insist that the sequence be increasing according to this notion. This naturally corresponds to the structure of an arbitrary tournament, recording for each pair $(a,b)$ whether $x_{a}<_{1}x_{b}$ or $x_{b}<_{1}x_{a}$.
@@
In this paper, we study a close cousin of these classical questions, which turns out to be surprisingly rich and difficult. On the level of vectors, we study a generalization of the relation $_{1}$ considered above, in which we ask whether one vector is strictly greater than another on at least $r$ of the coordinates. On the level of graph theory, we move from the classical Ramsey-theoretic goal of finding a monochromatic structure, to the alternative goal of finding a color-avoiding structure; that is, one which receives at most q1q-1 colors out of the palette of $q$ colors used in total. We now turn to discuss these problems, and our results, in more detail.
1.1. Sequences of vectors
Let $q\geqslant r\geqslant 1$ be integers. Extending the definition above, we define the relation $_{r}$ on $\mathbb{R}^{q}$ by saying that $x<_{r}y$ if there are at least $r$ coordinates $i\in[q]$ for which $x_{i}<y_{i}$. Despite the notation, $_{r}$ is not a partial order whenever $r<q$, as it is not transitive. There are (at least) two natural extremal questions one can ask about this relation, both of which have a long and rich history. First, following the notation of Gowers and Long [22], we define $F_{q,r}(n)$ to be the maximum number of vectors $x_{1},\dots,x_{N}\in[n]^{q}$ such that for all $a<b$, we have $x_{a}<_{r}x_{b}$. In other words, $F_{q,r}(n)$ measures the maximum length of an $r$-increasing sequence of vectors in $n]^{q}$. Second, we also study a weaker notion, of $r$-comparable sets of vectors; formally, we define $G_{q,r}(n)$ to be the maximum size of a set $\{x_{1},\dots,x_{N}\}\subseteq[n]^{q}$ such that for all $a\neq b$, either $x_{a}<_{r}x_{b}$ or $x_{b}<_{r}x_{a}$. Given such an $r$-comparable set, we naturally obtain an $N$-vertex tournament [^1] Strictly speaking, one only obtains a well-defined tournament if $r>q/2$, as this condition guarantees that if $x<_{r}y$, then $y\not<_{r}x$. If $r\leqslant q/2$, then we can, for example, allow antiparallel edges in case $x<_{r}y$ and $y<_{r}x$., recording for each pair $(a,b)$ whether $x_{a}<_{r}x_{b}$ or $x_{b}<_{r}x_{a}$. Thus, the $r$-comparable set is in fact an $r$-increasing sequence if and only if this tournament is transitive. Additionally, the fact that the $r$-comparable condition is weaker immediately implies that $F_{q,r}(n)\leqslant G_{q,r}(n)$ for all $q,r,n$.
The first non-trivial case of this problem is $(q,r)=(3,2)$, which has been extensively studied for over 50 years, with different researchers arriving at this question from a number of different perspectives, including extremal problems in convex geometry [6] and hypergraphs [22], $k$-majority tournaments [26], incidence geometry [36], the Erdős–Hajnal conjecture [39], and packing questions motivated by group theory [33] and error-correcting codes [24]. For a detailed discussion of these disparate instantiations of this problem, we refer to [3, Section 4.3]. For this problem, it is quite easy to observe the bounds
(1) $(1-o(1)){n^{3/2}}\leqslant F_{3,2}(n)\leqslant G_{3,2}(n)\leqslant n^{2}.$
The lower bound follows from an iterative product construction, which shows that $F_{3,2}(n)\geqslant n^{3/2}$ whenever $n$ is a perfect square (see e.g. [22, Proposition 1.1] for details). The upper bound follows from the simple observation that if we delete the last coordinate from a $2$-comparable set in $n]^{3}$, we obtain a $1$-comparable set in $n]^{2}$, which in particular must consist of distinct vectors. The same argument shows that for all integers $q\geqslant r>t$, we have
(2) $F_{q,r}(n)\leqslant F_{q-t,r-t}(n)\qquad\text{and}\qquad G_{q,r}(n)\leqslant G_{q-t,r-t}(n),$
since by deleting the last $t$ coordinates of an $r$-increasing sequence (resp. an $r$-comparable set) in $n]^{q}$, we obtain an (rt)(r-t)-increasing sequence (resp. an (rt)(r-t)-comparable set) in $n]^{q-t}$.
Improving the lower bound on $F_{3,2}(n)$ in (1) appears to be extremely difficult, and Gowers–Long [22, Conjecture 1.2] conjecture that in fact, $F_{3,2}(n)\leqslant n^{3/2}$ for all $n$. However, quite surprisingly, weakening the problem from $2$-increasing sequences in $n]^{3}$ to $2$-comparable subsets of $n]^{3}$ makes a big difference, and already in 1984 Hamaker and Stein [24] found a $2$-comparable subset of $n]^{3}$ of size $n^{1.51}$, implying that $G_{3,2}(n)\geqslant n^{1.51}$ for sufficiently large $n$. The exponent in the lower bound was steadily improved (e.g. [37, 34, 35]), always by finding a construction for some fixed $n$, and then using the same product argument to boost this construction to all large $n$. Unaware of these prior developments, Gowers and Long [22] developed a continuous relaxation of this problem, allowing them to improve the lower bound to $G_{3,2}(n)\geqslant n^{1.54}$; they even speculate, based on some computational evidence of Wagner, that their construction may be optimal.
On the upper bound side, the main question left open in [34, 24] is whether $G_{3,2}(n)=o(n^{2})$. This was finally proved by Tiskin [37], via a rather simple argument: if one views the elements of a $2$-comparable set of vectors in $n]^{3}$ as the edges of a tripartite $3$-uniform hypergraph on vertex set $n]\sqcup[n]\sqcup[n]$, then it is not hard to check that this hypergraph does not $6$ vertices spanning $3$ edges. By applying the famous $(6,3)$ theorem of Ruzsa and Szemerédi [30], one concludes that $G_{3,2}(n)=o(n^{2})$. However, the savings over the trivial upper bound are very minor; even by applying the strongest known quantitative version of the $(6,3)$ theorem, due to the first author [17], one only obtains that $G_{3,2}(n)\leqslant n^{2}/2^{\Omega(\log^{*}n)}$, where $\log^{*}n$ is the iterated logarithm function. To date, this remains the best known upper bound on $G_{3,2}(n)$.
The only other improvement on the trivial bounds in (1) to date is due to Gowers and Long, who proved that
(3) $F_{3,2}(n)\leqslant n^{2-\delta},$
@@ 1.1. Sequences of vectors
for an extremely small, but absolute, constant $\delta>0$. Their proof is based on an induction argument, and the base case of the induction is established by applying the bound $F_{3,2}(n)\leqslant G_{3,2}(n)\leqslant n^{2}/2^{\Omega(\log^{*}n)}$ discussed above; since this improvement only beats the trivial bound for an enormous $n$, the constant $\delta$ they obtain is unbelievably small.
For larger values of $q,r$, the simple bounds corresponding to (1) are
$(1-o(1))n^{q/r}\leqslant F_{q,r}(n)\leqslant G_{q,r}(n)\leqslant n^{q-r+1}.$
The lower bound again follows from the same product construction, and the upper bound follows by applying (2) with t=r1t=r-1, as well as the trivial bound $G_{q-r+1,1}(n)\leqslant n^{q-r+1}$. In general, the lower bound is far from tight; Gowers and Long [22, Lemma 5.1] noted that if $q$ is large and $r<(\frac{1}{2}-\gamma)q$ for some fixed $\gamma>0$, then a simple probabilistic argument gives a much better lower bound than $n^{q/r}$ for $n$ fixed; one can then use the same product construction to boost this lower bound to larger values of $n$. However, for $r>q/2$, as far as we are aware, the best known lower bound remains $F_{q,r}(n)\geqslant n^{q/r}$. For the special case r=q1r=q-1, these bounds become
(4) $(1-o(1))n^{q/(q-1)}\leqslant F_{q,q-1}(n)\leqslant G_{q,q-1}(n)\leqslant n^{2},$
and thus the gap in the exponent widens as $q$ grows. For upper bounds, one can apply the simple reductions (2) to obtain a power savings when r=q1r=q-1, as $F_{q,q-1}(n)\leqslant F_{3,2}(n)\leqslant n^{2-\delta}$, by the result of Gowers and Long [22]. However, no such power savings was known for the more general $r$-comparable question; the only upper bound, which follows from the hypergraph removal lemma as in Tiskin’s argument, is $G_{q,r}(n)=o(n^{q-r+1})$, where the savings is extraordinarily small (the savings is roughly the reciprocal of the (qr+1)(q-r+1)th level of the Ackermann hierarchy, evaluated at $n$). As such, Gowers and Long highlighted the following question, saying it is the “most annoying” question they were unable to answer.
Question 1.1 ([22, Question 6.3]).

### v2

# Color-avoiding directed paths in tournaments

@@ Abstract.
We study the following Ramsey-theoretic question: given a $q$-coloring of the edges of a tournament, how long of a directed path can we guarantee whose edges avoid one of the colors? Questions of this type have applications in many areas, such as vector sequences, convex geometry, and extremal hypergraph theory, and have been extensively studied over the past 50 years.
We prove that if $\varepsilon>0$ is fixed and $q$ is sufficiently large, then every $q$-edge-colored NN-vertex tournament contains a color-avoiding directed path of length $N^{1-\varepsilon}$. This answers a question of Gowers and Long, strengthens several of their results, and extends earlier work of Loh.
JF: Department of Mathematics, Stanford University, USA. Email: jacobfox@stanford.edu. Research supported by NSF awards DMS-2452737 and DMS-2154129
@@ 1. Introduction
1.1. Background and main results
Questions about directed paths in tournaments are as old as the study of tournaments itself; the very first result on tournaments is due to Rédei [32], who proved that every tournament contains a Hamiltonian directed path, i.e. a directed path visiting every vertex. A substantial generalization of this result was obtained independently by Gallai, Hasse, Roy, and Vitaver [24, 28, 33, 44], who proved that every edge-orientation of a graph $G$ contains a directed path of length at least $\chi(G)$; Rédei’s theorem then corresponds to the case where $G$ is a complete graph. Throughout this paper, by the length of a path, we mean the number of vertices it comprises.
Our focus in this paper will be on Ramsey-theoretic questions in tournaments, wherein we are given a tournament whose edges have been assigned one of $q$ possible colors, and we wish to understand which sorts of substructures must appear in this coloring. The study of such questions can be traced back to work of Stearns [37] and Erdős–Moser [18] from the late 1950s, and they have become increasingly well-studied in the past few years, e.g. [11, 22, 7, 46, 9].
One early result in this direction, due independently to Chvátal [12] and Gyárfás–Lehel [26], again concerns directed paths; they proved that in every $q$-edge-coloring of an $N$-vertex tournament, there must be a monochromatic directed path of length at least $N^{1/q}$. This fact is a simple corollary of the Gallai–Hasse–Roy–Vitaver theorem mentioned above: viewing each color class as an $N$-vertex graph, these $q$ graphs have $K_{N}$ as their union, hence one of them must have chromatic number at least $N^{1/q}$, from which the result follows. By using a simple product construction (see Proposition 3.2 for details), one can show that this bound is tight whenever $N$ is a power of $q$, even when restricting to transitive tournaments.
Despite its simple proof, even very special cases of the Chvátal–Gyárfás–Lehel theorem have a number of remarkable consequences. For example, restricting our attention to the case $q=2$, and focusing only on transitive tournaments, the result we obtain is equivalent to the famous Erdős–Szekeres lemma [19] on monotone subsequences, stating that every sequence of $N$ real numbers has a monotone subsequence of length at least $\sqrt{N}$. Indeed, any such sequence yields a two-coloring of a transitive $N$-vertex tournament, coloring an edge red if its right endpoint is greater than its left (and blue otherwise); it is easy to see that monotone subsequences correspond to monochromatic directed paths in this coloring, from which one deduces the Erdős–Szekeres lemma. Similarly, if we increase $q$ but still restrict to transitive tournaments, we readily recover the multidimensional analogue of the Erdős–Szekeres lemma. Further generalizations of these statements and arguments exhibit deep connections to discrete geometry [19] and high-dimensional partitions [30], among others.
@@ 1.1. Background and main results
In this paper, we study a close cousin of these classical questions, which turns out to be surprisingly rich and difficult. We move from the classical Ramsey-theoretic goal of finding a monochromatic structure, to the alternative goal of finding a color-avoiding structure; that is, one which receives at most q1q-1 colors out of the palette of $q$ colors used in total. The study of such questions goes back at least to the seminal work of Erdős and Szemerédi [20], and is a special case of the “generalized Ramsey theory” introduced by Erdős and Gyárfás [17] and intensively studied since, e.g. [31, 1, 10].
The most basic question of the kind we study was introduced by Loh [29], who asked for the length of the longest color-avoiding directed path that is guaranteed in every $3$-edge-coloring of an $N$-vertex tournament. We denote this quantity by $g_{3,2}(N)$. More generally, for integers $q\geqslant r\geqslant 1$, let us define $g_{q,r}(N)$ analogously: it is the minimum, over all $q$-edge-colored NN-vertex tournaments, of the number of vertices in the longest directed path whose edges receive at most $r$ colors. While our focus will be on the color-avoiding case, in which r=q1r=q-1, the general function is equally natural to study.
In introducing this question, Loh observed two simple upper and lower bounds on $g_{3,2}(N)$. First, we have the trivial lower bound $g_{3,2}(N)\geqslant\sqrt{N}$, obtained by simply merging two of the colors and applying the Chvátal–Gyárfás–Lehel theorem to find a monochromatic directed path in the modified coloring; any such path must receive at most two colors in the original coloring, and hence is color-avoiding. On the other hand, Loh noted that the same simple product construction witnessing the tightness of the Chvátal–Gyárfás–Lehel theorem can be used to show that $g_{3,2}(N)\leqslant N^{2/3}$ whenever $N$ is a perfect cube, and hence $g_{3,2}(N)\leqslant(1+o(1))N^{2/3}$ in general. In fact, the example witnessing this is a $3$-edge-colored transitive tournament; the case of transitive tournaments plays an important role (and was Loh’s main object of study), so it is natural to further define $f_{q,r}(N)$ identically to $g_{q,r}(N)$, except restricting our attention only to $q$-edge-colorings of the $N$-vertex transitive tournament. We thus have $g_{q,r}(N)\leqslant f_{q,r}(N)$ for all $q,r,N$, and the bounds discussed above show that
$\sqrt{N}\leqslant g_{3,2}(N)\leqslant f_{3,2}(N)\leqslant(1+o(1))N^{2/3}.$ (1)
It seems to be surprisingly difficult to meaningfully improve either of these bounds, and Loh was only able to show a very minor asymptotic improvement on the lower bound for $f_{3,2}(N)$, proving that $f_{3,2}(N)\geqslant\sqrt{N}\cdot 2^{\Omega(\log^{*}N)}$, where $\log^{*}$ is the iterated logarithm function. This is proved via a reduction to the famous triangle removal lemma of Ruzsa and Szemerédi [34], and the precise bound stated above follows from the strongest known quantitative version of the triangle removal lemma, due to the first author [21].
@@
Loh’s work inspired a number of further authors [25, 42, 31, 45] to study this kind of question. Notably, Gowers and Long [25] managed to significantly improve Loh’s lower bound for $f_{3,2}(N)$ by a polynomial amount, showing that $f_{3,2}(N)\geqslant N^{\frac{1}{2}+\delta}$ for an extremely small, but absolute, constant $\delta>0$. Their proof is based on an induction argument, and the base case of the induction is established by applying the bound $f_{3,2}(N)\geqslant\sqrt{N}\cdot 2^{\Omega(\log^{*}N)}$ discussed above; since this improvement only beats the trivial bound for an enormous $N$, the constant $\delta$ they obtain is tiny. To the best of our knowledge, there is no known asymptotic improvement on the lower bound $g_{3,2}(N)\geqslant\sqrt{N}$.
When it comes to upper bounds on $f_{3,2}(N)$, nothing better than $(1+o(1))N^{2/3}$ is known, and Gowers and Long [25] conjectured that $f_{3,2}(N)\geqslant N^{2/3}$. However, rather surprisingly, better upper bounds on $g_{3,2}(N)$ are available, and have been known (in a disguised form) since the 1980s. Indeed, Hamaker and Stein [27] found a construction for the closely related tripod packing problem (discussed further below) which, when converted back to our setting, yields a $3$-edge-colored NN-vertex tournament whose longest color-avoiding directed path has length $N^{0.663}$. The exponent in the upper bound has been steadily improved (e.g. [43, 40, 41]). Unaware of these prior developments, Gowers and Long [25] developed a continuous relaxation of this problem, allowing them to obtain current record of $g_{3,2}(n)\leqslant N^{0.649}$; they even speculate, based on some computational evidence of Wagner, that their construction may be optimal. We stress that all of these constructions use highly non-transitive tournaments, and that the best known upper bound in the transitive case remains $f_{3,2}(N)\leqslant(1+o(1))N^{2/3}$, which is conjecturally optimal; thus, in contrast to the Chvátal–Gyárfás–Lehel theorem, it seems that the transitive case is not extremal for this color-avoiding problem. This is a genuine difference from the monochromatic case, and may serve to explain some of the difficulty of this problem.
In this paper, we study this kind of color-avoiding question in $q$-colored tournaments, where $q\geqslant 4$. Extending (1), the simple known bounds on this problem are
$\sqrt{N}\leqslant g_{q,q-1}(N)\leqslant f_{q,q-1}(N)\leqslant(1+o(1))N^{1-1/q},$
where the upper bound follows from the same product construction and the lower bound again follows by arbitrarily merging the colors into two classes and applying the Chvátal–Gyárfás–Lehel theorem. Note that the gap between the upper and lower bounds only grows as $q$ increases; moreover, the lower bound appears to be more and more wasteful as $q$ grows. It is thus natural to expect that the true exponent on $N$ should increase with $q$, and perhaps even converge to $1$ as $q\to\infty$; this is precisely the statement of our main theorem.
Theorem 1.1.
We have that $g_{q,q-1}(N)\geqslant c_{q}N^{1-O(1/\sqrt{\log q})}$, where $c_{q}>0$ is a constant depending only on $q$. That is, in any $q$-edge coloring of any $N$-vertex tournament, there exists a color-avoiding directed path of length at least $c_{q}N^{1-O(1/\sqrt{\log q})}$.
In particular, we obtain the same lower bound when restricting to transitive tournaments; however, as we discuss below, alternative techniques give even stronger lower bounds on $f_{q,q-1}(N)$. Nevertheless, rather interestingly, our proof approach demonstrates that any extremal construction for $g_{q,q-1}(N)$ for $q\geqslant 4$ must be approximately transitive. To make this precise, let us recall that an $N$-vertex tournament is $\delta$-close to transitive if it can be made transitive by reversing the orientation of at most $\delta N^{2}$ edges, and is $\delta$-far from transitive otherwise.
Theorem 1.2.
If an $N$-vertex tournament is $\delta$-far from transitive, then any $q$-coloring of its edges contains a directed path of length $c\delta^{2}N/q^{3}$ which is colored by at most three colors, where $c>0$ is an absolute constant. Furthermore, the coloring contains the square of a path such that the direction of the edges between consecutive vertices is forward, the direction of edges between vertices of distance two is backwards, and the coloring of the edges is periodic with period $3$.
The proof of Theorem 1.2 is extremely short and simple, and relies on a result of the first two authors [23] on the structure of tournaments that are far from transitive. They proved a directed version of the triangle removal lemma [34] with very good quantitative bounds, and it is this good quantitative behavior that means that the $\delta$-dependence in Theorem 1.2 is polynomial.
1.2. Applications: vectors sequences and pod packings
All prior works studying color-avoiding directed paths have used a close connection to certain geometric questions, whose origin can be traced to Seidenberg’s classical proof [36] of the Erdős–Szekeres lemma. These geometric questions themselves have a long and storied history, and have been extensively studied for over 50 years, with different researchers arriving at them from a number of different perspectives, including extremal problems in convex geometry [8] and hypergraphs [25], $k$-majority tournaments [29], incidence geometry [42], the Erdős–Hajnal conjecture [45], and packing questions motivated by group theory [39] and error-correcting codes [27]. For a detailed discussion of these disparate instantiations, we refer to [4, Section 4.3].
To describe this connection, we make the following definition. Given integers $q\geqslant r\geqslant 1$, we define the relation $_{r}$ on $\mathbb{R}^{q}$ by saying that $x<_{r}y$ if there are at least $r$ coordinates $i\in[q]$ for which $x_{i}<y_{i}$. Despite the notation, $_{r}$ is not a partial order whenever $r<q$, as it is not transitive.
In order to explain how this relates to color-avoiding directed paths, let us focus for the moment on the transitive setting: say we are given a transitive tournament with vertex set $[N]$ and all edges oriented forwards, and suppose each of its edges has been assigned a color from $[q]$. For every vertex $a\in[N]$, we define a vector $x_{a}\in\mathbb{N}^{q}$ as follows: the $i$th coordinate of $x_{a}$ records the length of the longest directed path ending at $a$ whose edges avoid the color $i$. The property enjoyed by these vectors is that they form a (q1)(q-1)-increasing sequence. That is, for every $a<b$, we have that $x_{a}<_{q-1}x_{b}$. Indeed, if the color of the edge $ab$ is $i$, then for every $i\neq j\in[q]$, we may extend the longest $j$-avoiding path ending at $a$ to a strictly longer $j$-avoiding path ending at $b$, implying that $(x_{b})_{j}>(x_{a})_{j}$. Moreover, if $n$ is the length of the longest color-avoiding directed path in the given transitive tournament, then all the vectors $x_{a}$ lie in the box $n]^{q}$.
We are thus led to the natural extremal function $F_{q,q-1}(n)$, defined as the length of the longest (q1)(q-1)-increasing sequence of vectors in $n]^{q}$. The argument presented above shows if $f_{q,q-1}(N)\leqslant n$, then $F_{q,q-1}(n)\geqslant N$, as we may convert any transitive tournament with no long color-avoiding paths into a long (q1)(q-1)-increasing sequence. Somewhat surprisingly, this connection can be reversed: given a (q1)(q-1)-increasing sequence of $N$ vectors in $n]^{q}$, one can $q$-color the edges of an $N$-vertex transitive tournament such that every color-avoiding path has length at most $n$. Indeed, for $a<b$ one colors the edge $ab$ by the unique index $i$ (if it exists) such that $(x_{a})_{i}>(x_{b})_{i}$; if there is no such index, this edge can be colored arbitrarily. In this coloring, any directed path avoiding the color $i$ must have strictly increasing $i$th coordinates in the corresponding vectors, implying that its length is at most $n$ since the vectors are contained in $n]^{q}$. Combining these two constructions, we see that the extremal functions $F_{q,q-1}$ and $f_{q,q-1}$ are inverses of one another, in the sense that
$f_{q,q-1}(N)\leqslant n\qquad\text{if and only if}\qquad F_{q,q-1}(n)\geqslant N.$ (2)
For a formal proof of this fact, see e.g. [29, Lemma 2.1].
Of course, our study in this paper is not restricted to transitive tournaments, so it is natural to generalize the notion of (q1)(q-1)-increasing sequences as follows. We say that a set $\{x_{1},\dots,x_{N}\}\subseteq[n]^{q}$ is (q1)(q-1)-comparable if, for all $a<b$, we have $x_{a}<_{q-1}x_{b}$ or $x_{b}<_{q-1}x_{a}$, and we let $G_{q,q-1}(N)$ denote the maximum size of a (q1)(q-1)-comparable set of vectors in $n]^{q}$. Even more generally, for $q\geqslant r\geqslant 1$, we define $F_{q,r}(n)$ and $G_{q,r}(n)$ analogously, as the maximum length (resp. maximum size) of an $r$-increasing sequence (resp. $r$-comparable set) in $n]^{q}$. Given an $r$-comparable set, we naturally obtain an $N$-vertex tournament [^1] Strictly speaking, one only obtains a well-defined tournament if $r>q/2$, as this condition guarantees that if $x<_{r}y$, then $y\not<_{r}x$. If $r\leqslant q/2$, then we can, for example, allow antiparallel edges in case $x<_{r}y$ and $y<_{r}x$., recording for each pair $(a,b)$ whether $x_{a}<_{r}x_{b}$ or $x_{b}<_{r}x_{a}$. Thus, the $r$-comparable set is in fact an $r$-increasing sequence if and only if this tournament is transitive. Additionally, the fact that the $r$-comparable condition is weaker immediately implies that $F_{q,r}(n)\leqslant G_{q,r}(n)$ for all $q,r,n$.
Before proceeding, we make two warnings about the difficulty of extending (2). First, it is very natural to assume that this equivalence carries over verbatim to an equivalence between $F_{q,r}$ and $f_{q,r}$. However, this is not the case. Instead, the same argument above shows that $F_{q,r}(n)$ is equal to the largest $N$ for which the following holds. One can assign to every edge of an $N$-vertex transitive tournament a subset of $r$ colors from a palette of $q$ colors, and no color appears in the color set of every edge of a directed path $P$ of length $n+1$. This is one less than the set-coloring ordered Ramsey number of a monotone path with $n+1$ vertices. Set-coloring Ramsey numbers have been extensively studied recently for cliques (see [13, 15, 3]) and have found interesting connections to coding theory, and have also been studied for certain other classes of graphs (see e.g. [10] and the references therein). Ordered Ramsey numbers of graphs have also been extensively studied recently (see [14, 6] and the recent survey [5]).
Our second warning concerns the extension of (2) to the non-transitive setting. Again, it would be natural to expect that $G_{q,q-1}$ and $g_{q,q-1}$ are inverses of each other, but this is not the case. The thing that goes wrong is the first direction of the equivalence: it may be that the edge $a\to b$ is colored with color $i$, but we can nevertheless not extend a $j$-avoiding path ending at $a$ to one ending at $b$, as such a path may have already visited $b$. Note that we did not encounter this issue in the transitive setting, as such a situation can only arise in the presence of directed cycles. Although this issue may seem like a mere technicality, it turns out to be fundamental; it is also the main difficulty in the proof of the Gallai–Hasse–Roy–Vitaver theorem, and must also be carefully dealt with in our proof of Theorem 1.1. Nevertheless, the other direction of the equivalence does go through in the non-transitive setting: any (q1)(q-1)-comparable set of vectors in $n]^{q}$ yields a $q$-coloring of their corresponding tournament in which the longest directed color-avoiding path has at length at most $n$. That is, we have the one way implication
$\text{if }G_{q,q-1}(n)\geqslant N,\qquad\text{then }g_{q,q-1}(N)\leqslant n.$ (3)
Thus, for example, the constructions of Hamaker–Stein [27] and Gowers–Long [25] actually prove lower bounds on $G_{3,2}(n)$, from which one deduces the upper bounds on $g_{3,2}(N)$ which we discussed above.
In fact, the work of Hamaker–Stein, as well as many subsequent papers on this topic, study yet another geometric question which we now introduce. A tripod of order $n$ consists of a corner and the three adjacent edges of an integer $n\times n\times n$ cube. The corner of the tripod is called its apex. We say that a collection of tripods is aligned if they are translates of one other, and that they pack if they are pairwise disjoint. The maximum number of aligned tripods which pack and whose apex is in $n]^{3}$ is precisely $G_{3,2}(n)$, as two tripods with apices in $n]^{3}$ are disjoint if and only if their apices are $2$-comparable. Most of the early paper studying $G_{3,2}(n)$ were motivated from the question of packing tripods, and it is not hard to show that the two questions are equivalent: bounding the maximum number of disjoint tripods of order $n$ whose apex lies in $n]^{3}$ is equivalent to bounding the maximum density of $3$-space that can be packed by order-$n$ tripods. For a formal statement (and in particular for the formal definition of the density of a packing), we refer to [41, Chapter 4] or [38].
(a) A tripod of order 4
last coordinate = 1last coordinate = 2last coordinate = 0
Gowers and Long [25] proved a large number of results on the extremal functions $F_{q,r}(n)$ and $G_{q,r}(n)$. For example, we already mentioned their improved bound $f_{3,2}(N)\geqslant N^{\frac{1}{2}+\delta}$; they actually proved it in the equivalent form $F_{3,2}(n)\leqslant n^{2-\delta}$. It is easy to deduce from this that $F_{q,q-1}(n)\leqslant n^{2-\delta}$ for all $n$, again obtaining a power savings over the trivial bound of $n^{2}$. However, for the more general question of $G_{q,r}(n)$, their results were much more limited: the trivial upper bound here [^2] Indeed, by deleting the least r1r-1 coordinates from every vector in an $r$-comparable set, we obtain a $1$-comparable set in $n]^{q-r+1}$, whose size is trivially at most $n^{q-r+1}$. is $G_{q,r}(n)\leqslant n^{q-r+1}$, and they noted that an application of the hypergraph removal lemma, generalizing the use of the triangle removal lemma above [^3] In the setting of vectors, the fact that the removal lemma is applicable was first noted by Tiskin [43], several years before Loh [29] used the same idea in the study of color-avoiding paths., shows that $G_{q,r}(n)=o(n^{q-r+1})$. However, the savings is extraordinarily small (the savings is roughly the reciprocal of the (qr+1)(q-r+1)th level of the Ackermann hierarchy, evaluated at $n$). As such, Gowers and Long highlighted the following question, saying it is the “most annoying” question they were unable to answer.
Question 1.3 ([25, Question 6.3]).